the total no of valence electrons in 4.2g of n3-ion
 

Grade:12th pass

4 Answers

Aarti Gupta
9 years ago
In nitrogen(N) atom having atomic no.7,so electronic configuration = 2,5.
Thus valence e- = 5.Now when it forms N3- ion total no. of valence electrons = 8
Atomic mass of nitrogen = 14
so,
14g N3- ion have valence electrons = 8N (where N = Avogadro’s number)
So, 1g of N3- ion will have valence electrons = 8N/14
therefore,4.2g of N3-ion will have valence electrons = 8N* 4.2/14
= 2.4N
SO,4.2g of N3- ion will have 2.4N valence electrons.
Himanshu Malik
6 years ago
No of moles in 4.2g of N =4.2/14 = 3/10molesNow find out electrons in N^3- We know N have 7 electrons after taking 3e it will make N^3- hence total no. Of electrons is 7+3=10...... Now 4.2 gm N^3- i.e. 3/10 moles of this ion will have 3/10 ×10×NA(6.022 ×10^23) ions = 3NA electrons will be final answer
Himanshu Malik
6 years ago
In same way we can find valence electron ...total no. Of valence eletron = 5+3=8.. hence 3/10×8×NA = 2.4Na...........is the no. Of valence Electrons in given amount .
ankit singh
3 years ago
ion will have 2.4N valence electrons. 3- ion will have valence electrons = 8N* 4.2/14 = 2.4NSO,4.2g of N3- ion will have valence electrons = 8N/14therefore,4.2g of N3- ion have valence electrons = 8N (where N = Avogadro’s number)So, 1g of N3-14g N
 
No of moles in 4.2g of N =4.2/14 = 3/10molesNow find out electrons in N^3- We know N have 7 electrons after taking 3e it will make N^3- hence total no. Of electrons is 7+3=10...... Now 4.2 gm N^3- i.e. 3/10 moles of this ion will have 3/10 ×10×NA(6.022 ×10^23) ions = 3NA electrons will be final answer
 
In same way we can find valence electron ...total no. Of valence eletron = 5+3=8.. hence 3/10×8×NA = 2.4Na...........is the no. Of valence Electrons in given amount